# The frustrated Ising model on the honeycomb lattice: Metastability and universality

### v1

@@ IV.3 Finite-size scaling analysis
with $x$ and $\mathcal{O}$ being placeholders for a quantity $\mathcal{O}$ with the (asymptotic) FSS relation $\mathcal{O}\propto L^{x}$. Figure 10 shows the estimated slopes using the three-point midpoint approximation  [^3] The derivative is approximated using $x]_{\text{eff}}(\tilde{L}_{i})=\frac{\ln{O(L_{i+1})}-\ln{O(L_{i-1})}}{\ln L_{i+1}-\ln L_{i-1}},$ where $\tilde{L}_{i}=\exp[(\ln L_{i-1}+\ln L_{i+1})/2]=\sqrt{L_{i-1}L_{i+1}}$ is the geometric mean of $L_{i-1}$ and $L_{i+1}$. for Eq. (12). Despite the large statistical fluctuation of the numerical derivative, a clear variation of the effective exponents with $L$ can be seen for $\gamma/\nu$ and $1/\nu$. For the largest $L$ the slopes are always consistent with the expected asymptotic exponent. Note that the error bars of the effective exponents $\beta/\nu]_{\text{eff}}$ and $(1-\beta)/\nu]_{\text{eff}}$ are too large to precisely evaluate the relevance of scaling corrections.
@@ IV.4 Energy barriers
Finally, we consider the frustration-induced energy barriers present in this model, which form the building blocks of the rough free-energy landscape, and which turn out to provide a rather intuitive explanation of why our rejection-free approach outperforms the standard Metropolis algorithm for low $J_{2}$, and also why the approach fails for even lower $J_{2}$. The metastable states reported in Ref. [20] consist of multiple ferromagnetic domains with extremely long lifetimes. As $J_{2}$ is reduced the energetic barrier to overturn domains increases, and so does their lifetime. At the same time the energy difference between the ground state and a state with an overturned domain becomes smaller, rendering it less likely for the excited states to be filtered out during the resampling step in PA.
The simplest such domain on the honeycomb lattice corresponds to flipping a single hexagon. In fact, studying snapshots of low-temperature simulations reveals that indeed single hexagonal excitations happen frequently, and that they have a long lifetime; see the bottom panel of Fig. 11. For the following discussion we consider the difference in energy $\Delta E$ when overturning $k$ consecutive spins on the hexagons (see the top panel of Fig. 11). Higher excitations, e.g., branching or overturning non-consecutive spins on the hexagon are ignored, resulting in a very simple one-dimensional representation of the energy landscape.
Starting from the $k=6$ state (cf. the top panel of Fig. 11), the $n$-fold way update brings the system very efficiently to the $k=5$ state, as other available moves are typically even more expensive. Using regular Metropolis, the $6\rightarrow 5$ move depending on the temperature (which generally is lower the smaller $J_{2}$) might be rejected many times before eventually being accepted.
@@
Figure 11: Top: Difference in energy between the excited states and the ferromagnetic ground state for various $J_{2}$. $k$ denotes the number of overturned spins from the ground state. $k=6$ corresponds to one overturned hexagon (see illustration in the top right). For the calculation of $\Delta E$ we refer to the Appendix A and Table 2 therein. Bottom: Partial snapshot of an $L=64$ system for $J_{2}=-0.23$ and $\beta\approx 1.2\ \beta_{c}$.
@@
For $J_{2}<-1/5$ the energy difference $\Delta E(2\rightarrow 1)$ is larger than the one to go from $k=2$ to $k=3$, making the latter more likely to occur for the $n$-fold way update as well as for Metropolis. What happens is that the simulation oscillates between $k=2$ and $k=3$ for some time before reaching $k=1$. The further $J_{2}$ is reduced, the larger $\Delta E(2\rightarrow 1)$, and the smaller $\Delta E(2\rightarrow 3)$ becomes. Hence, as $J_{2}$ is reduced, reaching $k=1$ takes increasingly long. For example, at $T_{c}$ for $J_{2}=-0.23$ going from $k=2$ to $k=3$ is about 45 times more likely than going to $k=1$. Once reaching $k=1$ the chance of selecting $k=0$ next is $1/4$  [^4] Both $k=0$ and $k=2$ are energetically more favorable than $k=1$. As there are three nearest neighbors, there are three adjacent states corresponding to $k=2$, and only one state corresponding to $k=0$.. Thus, the system on average jumps between $k=1$, 22, and $3$ almost 200 times before reaching the ground state. This is why the $n$-fold way can still equilibrate the system for $J_{2}=-0.23$, but not as easily as for higher $J_{2}$. In contrast, for $J_{2}=-0.24$ assuming $T_{c}\approx 0.1$  [^5] Using exact enumeration (see Fig. 2) of a system with 32 spins, we find that the location of the specific-heat peak is at $0.1$. Thus, the actual critical temperature is well below $0.1$ and the $3000$ underestimates the number of spin flips necessary to reach the $k=1$ state. we find that the $k=2$ to $k=3$ transition is at least 3000 times more likely than $2\rightarrow 1$. Thus, ultimately, $n$-fold way PA cannot equilibrate $J_{2}=-0.24$ in any reasonable time. By deriving the transition matrix of this simplified model with seven states and considering its discrete phase-type distribution [61], one finds that it would take on average 500 spin flips ($2.3\times 10^{8}$ MCS) to go from $k=6$ to $k=0$ at $T_{c}$ for $J_{2}=-0.23$. For $J_{2}=-0.24$ at $T=0.1$, on the other hand, going from $k=6$ to $k=0$ would require 35 00035\,000 spin flips ($1.6\times 10^{13}$ MCS).
@@ V Summary and Outlook
In order to equilibrate our simulations we have replaced the standard Metropolis algorithm in the population annealing algorithm by the rejection-free $n$-fold way update [32] which allowed us to equilibrate lower $J_{2}$ than in previous work [20] where low acceptance rates led to quasi-frozen states. This shows, on the one hand, that the $n$-fold way update is not just a useful tool in non-equilibrium settings at low temperature but also well suited in equilibrium when no efficient cluster update is available. On the other hand, it highlights that population annealing is agnostic to the used update mechanism which previously was already realized in the context of molecular dynamics simulations [38]. This approach is further enhanced by introducing an adaptive number of sweeps into the population annealing framework, which works particularly well for the model at hand: Energy barriers cause simulations to be trapped in local energy minima at low temperatures. By using the adaptive sweep protocol, the population is quickly cooled at high temperatures when equilibration is easy and computing cost is focused on the temperature range where simulations tend to be trapped in these minima.
@@
Our method still fails to equilibrate at even smaller values for $J_{2}$, viz. $J_{2}\in(-0.25,-0.23)$. We attribute this to the fact that on the one hand the energies of configurations containing overturned domains are close to the ground-state energy, while on the other hand these configurations are separated from the ground state by large energetic barriers. This, on the one hand, makes it difficult for PA’s resampling to filter out configurations containing such domains, and, on the other, results in those domains having very long lifetimes. $J_{2}$ even closer to (but larger than) 0.25-0.25 could possibly be simulated by taking into account these excitations, for example by proposing to flip two neighboring sites (thus eliminating the $2\leftrightarrow 3$ oscillation in Fig. 11) or to flip entire hexagons. (An adaptation of the $n$-fold way algorithm to include updates of multiple spins simultaneously was proposed in Ref. [62].)
@@
Both, using the rejection-free update and the adaptive sweep schedule, as well as the combination thereof may be useful for studying other systems in which rugged free-energy landscapes and low acceptance rates render numerical simulations difficult. However, for our method to be suitable as a general-purpose approach, two issues require addressing. First, advancing all replicas to the maximum of the observed Metropolis times among them can deteriorate parallel efficiency when all replicas have to wait on one or a few other replicas. As the maximum is used, this problem becomes worse the larger the population size is. While this issue did affect our simulations to some extent, the extreme-value distribution of the Metropolis times in our case still allowed for efficient computation for the chosen population size. Second, the adaptive sweep protocol suggested here assumes that it is possible to decorrelate replicas through Monte Carlo updates. Clearly, when ergodicity is effectively broken, as is typically the case in glassy systems below the glass-transition temperature, the decorrelation criterion may not be satisfied within a reasonable time, and the simulation likely would fail to proceed  [^6] In fact, this is exactly the fate of our $J_{2}=-0.24$ simulations.. In such cases the recently proposed adaptive sweep schedule [44] using the “restricted autocorrelation time” may be more suitable.
@@ Acknowledgements.
This project was supported by the Deutsch-Französische Hochschule (DFH-UFA) through the Doctoral College $\mathbb{L}^{4}$” under Grant No. CDFA-02-07. We further acknowledge support by the Leipzig Graduate School of Natural Sciences “BuildMoNa”. Most calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium.
Appendix A Energies for six states leading to the hexagonal excitation
@@
Figure 12: Domain decomposition into four subsets of non-interacting spin sites for the $J_{1}$-$J_{2}$ Ising model on the honeycomb lattice of linear system size $L=4$ containing $N=2L^{2}=32$ sites. Each color refers to a different subset. (a) Representation in “real” coordinates, and (b) in “memory” coordinates, showing how spins are arranged in a two-dimensional array structure.
Appendix B GPU domain decomposition
Since the sites are arranged regularly in the considered honeycomb model, and since each spin only interacts with a small number of spins within its vicinity, a domain decomposition in the spirit of the well-known checkerboard decomposition [64] of the square lattice with even side lengths $L$ is possible. Due to the nn and nnn interactions, it is not possible to decompose the lattice into two sublattices of non-interacting sites. Instead, four colors are required, but otherwise the procedure is analogous to the one of the checkerboard decomposition; see Fig. 12(a). In memory, the spins are represented in a two-dimensional array: In Fig. 12(b) the same decomposition is displayed in “memory” coordinates. Using this decomposition, one sweep consists of updating i) all blue sites, ii) all red sites, iii) all olive sites, and iv) all black sites (or any other order thereof), where in each step $N/4$ spins are updated in parallel.

### v2

@@ IV.3 Finite-size scaling analysis
with $x$ and $\mathcal{O}$ being placeholders for a quantity $\mathcal{O}$ with the (asymptotic) FSS relation $\mathcal{O}\propto L^{x}$. Figure 11 shows the estimated slopes using the three-point midpoint approximation  [^3] The derivative is approximated using $x]_{\text{eff}}(\tilde{L}_{i})=\frac{\ln{O(L_{i+1})}-\ln{O(L_{i-1})}}{\ln L_{i+1}-\ln L_{i-1}},$ where $\tilde{L}_{i}=\exp[(\ln L_{i-1}+\ln L_{i+1})/2]=\sqrt{L_{i-1}L_{i+1}}$ is the geometric mean of $L_{i-1}$ and $L_{i+1}$. for Eq. (13). Despite the large statistical fluctuation of the numerical derivative, a clear variation of the effective exponents with $L$ can be seen for $\gamma/\nu$ and $1/\nu$ (obtained using the peak locations of $\mathrm{d}\hskip-0.50003pt\ln|m|$). For the largest $L$ the slopes are always consistent with the expected asymptotic exponent. Note that the error bars of the effective exponents $\beta/\nu]_{\text{eff}}$ and $(1-\beta)/\nu]_{\text{eff}}$ are too large to precisely evaluate the relevance of scaling corrections.
@@ IV.4 Energy barriers
Finally, we consider the frustration-induced energy barriers present in this model, which form the building blocks of the rough free-energy landscape, and which turn out to provide a rather intuitive explanation of why our rejection-free approach outperforms the standard Metropolis algorithm for low $J_{2}$, and also why the approach fails for even lower $J_{2}$. The metastable states reported in Ref. [20] consist of multiple ferromagnetic domains with extremely long lifetimes. As $J_{2}$ is reduced the energetic barrier to overturn domains increases as does their lifetime. At the same time the energy difference between the ground state and a state with an overturned domain becomes smaller, rendering it less likely for the excited states to be filtered out during the resampling step in PA.
The simplest such domain on the honeycomb lattice corresponds to flipping a single hexagon. In fact, studying snapshots of low-temperature simulations reveals that indeed single hexagonal excitations happen frequently, and that they have a long lifetime; see the bottom panel of Fig. 12. For the following discussion we consider the difference in energy $\Delta E$ when overturning $k$ consecutive spins on the hexagons (see the top panel of Fig. 12). Higher excitations, e.g., branching or overturning nonconsecutive spins on the hexagon are ignored, resulting in a very simple one-dimensional representation of the energy landscape.
Starting from the $k=6$ state (cf. the top panel of Fig. 12), the $n$-fold way update brings the system very efficiently to the $k=5$ state, as other available moves are typically even more expensive. Using regular Metropolis, the $6\rightarrow 5$ move depending on the temperature (which generally is lower the smaller $J_{2}$) might be rejected many times before eventually being accepted. Note that the jump in energy from $k=6$ to $k=5$ increases with decreasing $J_{2}$, in line with the general statement about barriers and lifetimes provided above.
@@
Figure 12: Top: Difference in energy between the excited states and the ferromagnetic ground state for various $J_{2}$. $k$ denotes the number of overturned spins from the ground state. $k=6$ corresponds to one overturned hexagon (see illustration in the top right). For the calculation of $\Delta E$ we refer to Appendix C and Table 2 therein. Bottom: Partial snapshot of an $L=64$ system for $J_{2}=-0.23$ and $\beta\approx 1.2\ \beta_{c}$.
@@
For $J_{2}<-1/5$ the energy difference $\Delta E(2\rightarrow 1)$ is larger than the one to go from $k=2$ to $k=3$, making the latter more likely to occur for the $n$-fold way update as well as for Metropolis. What happens is that the simulation oscillates between $k=2$ and $k=3$ for some time before reaching $k=1$. The further $J_{2}$ is reduced, the larger $\Delta E(2\rightarrow 1)$, and the smaller $\Delta E(2\rightarrow 3)$ becomes. Hence, as $J_{2}$ is reduced, reaching $k=1$ takes increasingly long. For example, at $T_{c}$ for $J_{2}=-0.23$ going from $k=2$ to $k=3$ is about 45 times more likely than going to $k=1$. Once reaching $k=1$ the chance of selecting $k=0$ next is $1/4$  [^4] Both $k=0$ and $k=2$ are energetically more favorable than $k=1$. As there are three nearest neighbors, there are three adjacent states corresponding to $k=2$, and only one state corresponding to $k=0$.. Thus, the system on average jumps between $k=1$, 22, and $3$ almost 200 times before reaching the ground state. This is why the $n$-fold way can still equilibrate the system for $J_{2}=-0.23$, but not as easily as for higher $J_{2}$. In contrast, for $J_{2}=-0.24$ assuming $T_{c}\approx 0.1$  [^5] Using exact enumeration (see Fig. 2) of a system with 32 spins, we find that the location of the specific-heat peak is at $0.1$. Thus, the actual critical temperature is well below $0.1$ and the $3000$ underestimates the number of spin flips necessary to reach the $k=1$ state. we find that the $k=2$ to $k=3$ transition is at least 3000 times more likely than $2\rightarrow 1$. Thus, ultimately, the $n$-fold way PA cannot equilibrate $J_{2}=-0.24$ in any reasonable time. By deriving the transition matrix of this simplified model with seven states and considering its discrete phase-type distribution [61], one finds that it would take on average 500 spin flips ($2.3\times 10^{8}$ MCS) to go from $k=6$ to $k=0$ at $T_{c}$ for $J_{2}=-0.23$. For $J_{2}=-0.24$ at $T=0.1$, on the other hand, going from $k=6$ to $k=0$ would require 35 00035\,000 spin flips ($1.6\times 10^{13}$ MCS).
@@ V Summary and Outlook
In order to equilibrate our simulations we have replaced the standard Metropolis algorithm in the population annealing algorithm by the rejection-free $n$-fold way update [32], which allowed us to equilibrate lower $J_{2}$ than in previous work [20] where low acceptance rates led to quasifrozen states. This shows, on the one hand, that the $n$-fold way update is not just a useful tool in nonequilibrium settings at low temperature but also well suited in equilibrium when no efficient cluster update is available. On the other hand, it highlights that population annealing is agnostic to the used update mechanism which previously was already realized in the context of molecular dynamics simulations [38]. This approach is further enhanced by introducing an adaptive number of sweeps into the population annealing framework, which works particularly well for the model at hand: Energy barriers cause simulations to be trapped in local energy minima at low temperatures. By using the adaptive sweep protocol, the population is quickly cooled at high temperatures when equilibration is easy and computing cost is focused on the temperature range where simulations tend to be trapped in these minima.
@@
Our method still fails to equilibrate at even smaller values for $J_{2}$, viz. $J_{2}\in(-0.25,-0.23)$. We attribute this to the fact that the energies of configurations containing overturned domains are close to the ground-state energy and that these configurations are separated from the ground state by large energetic barriers. This, on the one hand, makes it difficult for PA’s resampling to filter out configurations containing such domains, and, on the other, results in those domains having very long lifetimes. $J_{2}$ even closer to (but larger than) 0.25-0.25 could possibly be simulated by taking into account these excitations, for example by proposing to flip two neighboring sites (thus eliminating the $2\leftrightarrow 3$ oscillation in Fig. 12) or to flip entire hexagons. (An adaptation of the $n$-fold way algorithm to include updates of multiple spins simultaneously was proposed in Ref. [62].)
@@
Both, using the rejection-free update and the adaptive sweep schedule, as well as the combination thereof may be useful for studying other systems in which rugged free-energy landscapes and low acceptance rates render numerical simulations difficult. However, for our method to be suitable as a general-purpose approach, two issues require addressing. First, advancing all replicas to the maximum of the observed Metropolis times among them can deteriorate parallel efficiency when all replicas have to wait for one or a few other replicas. As the maximum is used, this problem becomes worse the larger the population size is. While this issue did affect our simulations to some extent, the extreme-value distribution of the Metropolis times in our case still allowed for efficient computation for the chosen population size. Second, the adaptive sweep protocol suggested here assumes that it is possible to decorrelate replicas through Monte Carlo updates. Clearly, when ergodicity is effectively broken, as is typically the case in glassy systems below the glass-transition temperature, the decorrelation criterion may not be satisfied within a reasonable time, and the simulation likely would fail to proceed  [^6] In fact, this is exactly the fate of our $J_{2}=-0.24$ simulations.. In such cases the recently proposed adaptive sweep schedule [44] using the “restricted autocorrelation time” may be more suitable.
Note added. After the submission of this paper we became aware of a related Wang-Landau simulation study by Azhari et al. [64], which also investigates the critical behavior of this model.
@@ Acknowledgements.
This project was supported by the Deutsch-Französische Hochschule (DFH-UFA) through the Doctoral College $\mathbb{L}^{4}$” under Grant No. CDFA-02-07. We further acknowledge support by the Leipzig Graduate School of Natural Sciences “BuildMoNa”. M.W. acknowledges support from Emory University for a sabbatical stay, where part of this work was completed. Most calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant No. EP/T022108/1 and the HPC Midlands+ consortium.
Figure 13: Domain decomposition into four subsets of noninteracting spin sites for the $J_{1}$-$J_{2}$ Ising model on the honeycomb lattice of linear system size $L=4$ containing $N=2L^{2}=32$ sites. Each color refers to a different subset. (a) Representation in “real” coordinates, and (b) in “memory” coordinates, showing how spins are arranged in a two-dimensional array structure $L_{x}\times L_{y}$ with $L_{x}=2L$ and $L_{y}=L$.
Appendix A GPU domain decomposition
Since the sites are arranged regularly in the considered honeycomb model, and since each spin only interacts with a small number of spins within its vicinity, a domain decomposition in the spirit of the well-known checkerboard decomposition [65] of the square lattice with even side lengths $L$ is possible. Due to the NN and NNN interactions, it is not possible to decompose the lattice into two sublattices of noninteracting sites. Instead, four colors are required, but otherwise the procedure is analogous to the one of the checkerboard decomposition; see Fig. 13(a). In memory, the spins are represented in a two-dimensional array: In Fig. 13(b) the same decomposition is displayed in “memory” coordinates. Using this decomposition, one sweep consists of updating i) all blue sites, ii) all red sites, iii) all olive sites, and iv) all black sites (or any other order thereof), where in each step $N/4$ spins are updated in parallel.
Appendix B Binder parameter crossings
In contrast to the critical exponents, $U^{*}$ is only weakly universal, i.e., it depends on details of the lattice structure. The numerical estimates for $U^{*}$ quoted in the main text were produced as follows. $U_{4}(T_{c},L)$ was determined for the different system sizes $L$ at the previously obtained values of $T_{c}$ using (single) histogram reweighting from the closest simulation point; see the top panel of Fig. 14. Besides statistical fluctuations in $U_{4}$ and (unknown) finite-size corrections affecting the estimates for $U^{*}$, the inaccuracy of $T_{c}$ results in an additional systematic error contribution. Assuming that the two contributions are independent, the displayed error bars were obtained as the square root of the sum of the squared error contributions, that is, for example, the error bar of $0.0012$ for $J_{2}=-0.21$ and $L=64$ contains a contribution from the statistical error of $0.00080$ and one from the systematic error of $0.00081$. As can be seen in the top panel of Fig. 14, for small $L$ the $U_{4}$ values differ significantly from the rest due to finite-size corrections, whereas for larger systems the estimate is affected by the increasing total error. Thus, the system size at which $U^{*}$ is determined has to be selected somewhat carefully.
Due to how the simulations are set up, the statistical error is roughly system-size independent, whereas the systematic error scales with the slope of $U_{4}$ at $T_{c}$, i.e., with $L^{1/\nu}=L$, which is the reason for the error bars increasing with the system size in the top panel. This effect can be seen in the bottom panel of Fig. 14. The quoted values for $U^{*}$ in the main text, i.e., $0.6124(12)$, $0.6120(11)$, and $0.6117(11)$ for $J_{2}=-0.21$, 0.22-0.22, and 0.23-0.23, were obtained using $U_{4}(T_{c},L)$ of the smallest $L$ for which the systematic error due to the uncertainty in $T_{c}$ is greater than the statistical error. (The data points used for the final estimates for $U^{*}$ are highlighted in the top panel by large crosses.) For $J_{2}=-0.21$ and 0.22-0.22 this $L$ is 64, and for $J_{2}=-0.23$ it is 24 (the difference being due to the different accuracy in $T_{c}$). Note, that these values are within error bars of the value of the pure Ising model on the triangular lattice, $U^{*}=0.611827739(14)$ (Ref. [66]), but in any case even the (different) value $U^{*}=0.6106901(14)$ of the square lattice [66] is rather close.
Figure 14: Binder parameter $U_{4}$ at $T_{c}$ for different $J_{2}$ and $L$. Top: $U_{4}$ vs $1/L$ at $T_{c}$. The error bars correspond to the square root of the sum of the squared systematic and statistical errors. The values quoted in the text are highlighted by large crosses ($\times$). As reference values the exact values [66] of the triangular and the square lattice are indicated by colored lines. Bottom: The statistical error (dashed lines) and the systematic error (solid lines) as a function of $L$ for different $J_{2}$ (indicated by different colors).
Appendix C Energies for six states leading to the hexagonal excitation